write the slope-intercept form of the equation of each line
3x-2y= -16

Answer by Alan3354(6097)

write the slope-intercept form of the equation of each line
3x-2y= -16
-----------
That means solve for y
3x-2y= -16
2y = 3x + 16
y = (3/2)x + 8